# Can Pre-training Indicators Reliably Predict Fine-tuning Outcomes of LLMs?

### v1

@@ Abstract
While metrics available during pre-training, such as perplexity, correlates well with model performance at scaling-laws studies, their predictive capacities at a fixed model size remains unclear, hindering effective model selection and development. To address this gap, we formulate the task of selecting pre-training checkpoints to maximize downstream fine-tuning performance as a pairwise classification problem: predicting which of two LLMs, differing in their pre-training, will perform better after supervised fine-tuning (SFT). We construct a dataset using 50 1B parameter LLM variants with systematically varied pre-training configurations, e.g., objectives or data, and evaluate them on diverse downstream tasks after SFT. We first conduct a study and demonstrate that the conventional perplexity is a misleading indicator. As such, we introduce novel unsupervised and supervised proxy metrics derived from pre-training that successfully reduce the relative performance prediction error rate by over 50%. Despite the inherent complexity of this task, we demonstrate the practical utility of our proposed proxies in specific scenarios, paving the way for more efficient design of pre-training schemes optimized for various downstream tasks.
@@ 2.1 LLM Variants and Target SFT Tasks
To approximate pre-training studies while maintaining reasonable computational resources, we continuously trained a 1B parameter LLM with 100B tokens, systematically ablating pre-training objectives, data mixture re-weighting, and data filtering and tagging. This continuous pre-training approach allowed us to generate a wider range of model variants while managing computational resources. Pre-training objectives: We explored seven pre-training objectives: causal language modeling (CLM) (Brown et al., 2020), span corruption (SC) (Raffel et al., 2020), prefix language modeling (PLM) (Raffel et al., 2020), SC+CLM, UL2 (Tay et al., 2023a), UL2R (Tay et al., 2023b), and UL2R+CLM (Garcia et al., 2023). CLM and PLM generate tokens left-to-right, with CLM using the full context and PLM conditioning on a prefix. SC reconstructs masked spans, parameterized by noise density and mean span length, set to $(0.15,3)$ following (Raffel et al., 2020). SC+CLM jointly trains SC and CLM. UL2 mixes six SC variants with PLM, while UL2R uses two SC settings—$(0.15,3)$ and $(0.5,32)$—with PLM. UL2R+CLM extends UL2R by adding a CLM objective. Mixture re-weighting: We train on the 627B-token Slimpajama corpus (Soboleva et al., 2023), which includes seven diverse domains. We reweigh different domains following  (Shen et al., 2024), producing six 100B-token subsets by adjusting domain distributions (detailed in Table 3 in Appendix); Data filtering and tagging: Source domain metadata was integrated by pre-pending each instance with its respective domain label (e.g., [Common Crawl]). Length-based sub-corpora were generated by selecting instances within the [25%, 75%] and [75%, 100%] token length quantiles. We in total produced 50 distinct LLM variants, the specifications of which are provided in Table  4 in Appendix.
@@ 3 Predictive Power on SFT Tasks
Table 1 details the pairwise SFT prediction accuracy of various proxy metrics across 50 LLM variants. Conventional perplexity (PPL-CLM) exhibited low accuracy (e.g., 0.3 on SFT-CMS), contrasting sharply with its known correlation strength in scaling studies. The span corruption perplexity (PPL-SC) performed better ($>$ 0.5 accuracy), consistent with prior findings on span corruption benefits (UL2) (Tay et al., 2023a). Few-shot (k-shot) proxies achieved higher accuracy still, with Kshot-RAG reaching $\approx$ 0.7 on SFT-CMS and SFT-RAG. Despite these improvements, no single proxy proved universally reliable across all tested SFT tasks.
@@
To understand proxy limitations, we analyzed how well PPL-CLM, PPL-SC, and Kshot-RAG predict relative SFT performance between models differing only in their pre-training objective. We grouped models by objective (CLM, SC, UL2, etc.) and evaluated pairwise prediction accuracy for comparisons between these groups (details in Figure 2; Appendix B covers data variations). Confirming earlier results, PPL-SC and Kshot-RAG consistently outperformed PPL-CLM. However, their accuracy depended significantly on two factors: (1) The specific pre-training difference: Proxies better captured large performance gaps caused by different objectives (e.g., SC vs. CLM, often $\geq 0.6$ accuracy) than smaller variations. (2) The target SFT task: A specific comparison (e.g., SC vs. SC+CLM) could yield low accuracy on one task (SFT-CMS, 0.2) but high accuracy on others (SFT-RAG/SFT-CBQA, $\geq 0.6$).
@@ 4 Learning to Compare
Kshot-RAG .687$\pm$.073 .724$\pm$.047 .683$\pm$.077
Combine Five Proxies .612$\pm$.055 .585$\pm$.051 .540$\pm$.104
Learning To Compare (Relative to Kshot-RAG)
@@
4.1 Formulation
[…]
4.2 Experiment Setup
We implemented the supervised classifier using LightGBM (details on other models in Appendix Section C), training separate models per SFT task (CMS, RAG, CBQA). To ensure robustness, we performed 20 runs, each using a random 60%/40% split of the 50 LLM variants to generate training/testing pairs (splits varied per run). We report mean accuracy and standard deviation over the 20 runs in Table 2 (middle section), compared against unsupervised baselines including PPL-CLM and Kshot-RAG.
4.3 Results
@@ 5.2 Recall the Best Model from a Small Candidate Set
One key practical use for LLM performance predictors is to identify the most promising models within a group of candidates, which can lead to significant cost savings by reducing the number of models that undergo supervised fine-tuning. To assess our classifier’s effectiveness in this critical application—specifically, its ability to recall the best pre-trained LLMs—we performed a ranking experiment where pairwise comparisons between models were predicted and then aggregated into an overall ranking using Borda Count scoring (detailed in Appendix DDwork et al. (2001). Models achieving more pairwise ’wins’ received higher scores, indicating better predicted performance. The evaluation results, presented as top-1 and top-5 recall in Figure 5, show that our “learning-to-compare” method consistently identified the top-performing LLMs. Impressively, it achieved perfect top-1 recall for the SFT-CMS, SFT-RAG, and SFT-CBQA tasks by focusing on the top 7, 7, and 8 predicted models respectively, demonstrating its effectiveness even when narrowing down a relatively small candidate pool (as few as 8 models). Additionally, the unsupervised Kshot-RAG method showed strong performance, corroborating observations from Section 3.
@@ 6 Related Work
LLM pre-training fundamentally shapes capabilities like reasoning (Wei et al., 2022; Kojima et al., 2022; Zellers et al., 2019), knowledge (Chang et al., 2024), and tool use (Yao et al., 2023; Mo et al., 2023). Critical pre-training design choices include the training objective—such as dominant CLM (Brown et al., 2020; OpenAI, 2023) for generation, SC (Raffel et al., 2020) which aids fine-tuning (Tay et al., 2023a), or combined UL2-style approaches (Tay et al., 2023a, b; Garcia et al., 2023) potentially using PrefixLM (Du et al., 2022; Chowdhery et al., 2023)—and pre-trained corpus composition, which involves quality curation (Rae et al., 2021; Touvron et al., 2023), filtering (Penedo et al., 2023; Xia et al., 2024), and source mixing (Weber et al., 2024; Shen et al., 2024) to ensure broad coverage and robustness. Given the variety of design options, lightweight methods to predict final performance are highly desirable for efficient model development. This work investigates predictors for supervised fine-tuning outcomes, utilizing systematic variations across several pre-training design factors in our study.
The ability to predict the performance of large language models (LLMs) after fine-tuning has gained significant importance, largely driven by the substantial computational investment required for pre-training. Previous research (Kaplan et al., 2020; Hoffmann et al., 2022; Henighan et al., 2020) established scaling laws showing that increasing pre-training FLOPs typically reduces perplexity on held-out data, correlating with enhancements in capabilities like chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022), preference alignment (Ouyang et al., 2022; Bai et al., 2022), and multilingual understanding (Chowdhery et al., 2023), suggesting larger models generally yield better downstream performance. Analogous scaling phenomena, where lower perplexity often corresponds to improved outcomes, have also been noted when fine-tuning LLMs for specific applications (Zhang et al., 2024; Isik et al., 2025); for instance, Isik et al. (2025) reported such a correlation for machine translation performance. Nevertheless, the dependability of perplexity as a universal predictor has recently come under scrutiny in certain contexts, particularly for tasks involving long-context generation (Liu et al., 2024b) or many-shot in-context learning (Agarwal et al., 2024), implying it may not be a robust indicator across all downstream tasks.
@@ 7 Conclusion and Future Directions
Future research could focus on validating the generalizability of these methods across larger model scales, a wider range of tasks, and different fine-tuning techniques. Furthermore, investigating whether signals from intermediate checkpoints during long pre-training cycles can predict final fine-tuning outcomes represents an important research topic not covered here, especially for very large language models.
@@ Acknowledgements
Appendix A Pretraining and LLMs
@@ 35 UL2R+CLM DC-0 1e-3
Appendix B Proxy Predictive Accuracy
@@
Appendix C Classifier Implementation Detail
@@
The implementation details are as follows: For logistic regression, we use scikit-learn’s (Pedregosa et al., 2011) LogisticRegression with the default lbfgs solver for binary classification. The model applies $L_{2}$ regularization with strength $C=1.0$, fits an intercept, and runs up to 100 iterations. Class weighting is not applied. For the neural network, we use scikit-learn’s MLPClassifier with two hidden layers of size 32 each and ReLU activation. The model is optimized using the Adam solver and trained for a maximum of 100 iterations. All other hyperparameters are set to their default values. For LightGBM, we use the LGBMClassifie from the official lightgbm library 111 https://lightgbm.readthedocs.io/en/latest/pythonapi/lightgbm.LGBMClassifier.html. The objective is set to binary with binary_logloss as the evaluation metric. All other hyperparameters follow the default settings: num_leaves=31, learning_rate=0.1, n_estimators=100, feature_fraction=1.0, bagging_fraction=1.0, and no regularization (lambda_l1=0.0, lambda_l2=0.0).
Appendix D Ranking using Borda Count
We adopt a Borda Count-style scoring method Dwork et al. (2001) to transform the pairwise prediction between models to a global ranking. For each model $m_{i}$, we compute its total score by counting the number of pairwise wins over all other models.
$\text{Score}(m_{i})=\sum_{j\neq i}\mathbb{1}\left(f(m_{i},m_{j})>0.5\right),$
where $f(m_{i},m_{j})$ denotes the classifier’s predicted probability that $m_{i}$ outperforms $m_{j}$. $\mathbb{1}(\cdot)$ is the indicator function. Finally, models are ranked based on their total scores, with higher scores indicating better predicted fine-tuned performance.

### v2

# When Scale is Fixed: Revisiting Pre-training Indicators for LLM Fine-tuning Performance

@@ Abstract
While pre-trained available metrics, such as perplexity, correlates well with model performance at scaling-laws studies, their predictive capacities at a fixed model size remains unclear, hindering effective model selection and development. To address this gap, we formulate the task of selecting pretraining checkpoints to maximize downstream fine-tuning performance as a pairwise classification problem: predicting which of two LLMs, differing in their pre-training, will perform better after supervised fine-tuning (SFT). We construct a dataset using 50 1B parameter LLM variants with systematically varied pre-training configurations, e.g., objectives or data, and evaluate them on diverse downstream tasks after SFT. We first conduct a study and demonstrate that the conventional perplexity is a misleading indicator. As such, we introduce novel unsupervised and supervised proxy metrics derived from pre-training that successfully reduce the relative performance prediction error rate by over 50%. Despite the inherent complexity of this task, we demonstrate the practical utility of our proposed proxies in specific scenarios, paving the way for more efficient design of pre-training schemes optimized for various downstream tasks.
@@ 2.1 LLM Variants and Target SFT Tasks
To approximate pre-training studies while maintaining reasonable computational resources, we continuously trained a 1B parameter LLM with 100B tokens, systematically ablating pre-training objectives, data mixture re-weighting, and data filtering and tagging. This continuous pre-training approach allowed us to generate a wider range of model variants while managing computational resources. Pre-training objectives: We explored seven pre-training objectives: causal language modeling (CLM) (Brown et al., 2020), span corruption (SC) (Raffel et al., 2020), prefix language modeling (PLM) (Raffel et al., 2020), SC+CLM, UL2 (Tay et al., 2023a), UL2R (Tay et al., 2023b), and UL2R+CLM (Garcia et al., 2023). CLM and PLM generate tokens left-to-right, with CLM using the full context and PLM conditioning on a prefix. SC reconstructs masked spans, parameterized by noise density and mean span length, set to $(0.15,3)$ following (Raffel et al., 2020). SC+CLM jointly trains SC and CLM. UL2 mixes six SC variants with PLM, while UL2R uses two SC settings—$(0.15,3)$ and (0.5,32)(0.5,32)—with PLM. UL2R+CLM extends UL2R by adding a CLM objective. Mixture re-weighting: We train on the 627B-token Slimpajama corpus (Soboleva et al., 2023), which includes seven diverse domains. We reweigh different domains following (Shen et al., 2024), producing six 100B-token subsets by adjusting domain distributions (detailed in Table 3 in Appendix); Data filtering and tagging: Source domain metadata was integrated by pre-pending each instance with its respective domain label (e.g., [Common Crawl]). Length-based sub-corpora were generated by selecting instances within the [25%, 75%] and [75%, 100%] token length quantiles. We in total produced 50 distinct LLM variants, the specifications of which are provided in Table 4 in Appendix.
@@ 3 Predictive Power on SFT Tasks
Table 1 details the pairwise SFT prediction accuracy of various proxy metrics across 50 LLM variants. Conventional perplexity (PPL-CLM) exhibited low accuracy (e.g., 0.3 on SFT-CMS), contrasting sharply with its known correlation strength in scaling studies. The span corruption perplexity (PPL-SC) performed better (>> 0.5 accuracy), consistent with prior findings on span corruption benefits (UL2) (Tay et al., 2023a). Few-shot (k-shot) proxies achieved higher accuracy still, with Kshot-RAG reaching $\approx$ 0.7 on SFT-CMS and SFT-RAG. Despite these improvements, no single proxy proved universally reliable across all tested SFT tasks.
@@
To understand proxy limitations, we analyzed how well PPL-CLM, PPL-SC, and Kshot-RAG predict relative SFT performance between models differing only in their pre-training objective. We grouped models by objective (CLM, SC, UL2, etc.) and evaluated pairwise prediction accuracy for comparisons between these groups (details in Figure 2; Appendix C covers data variations). Confirming earlier results, PPL-SC and Kshot-RAG consistently outperformed PPL-CLM. However, their accuracy depended significantly on two factors: (1) The specific pre-training difference: Proxies better captured large performance gaps caused by different objectives (e.g., SC vs. CLM, often $\geq 0.6$ accuracy) than smaller variations. (2) The target SFT task: A specific comparison (e.g., SC vs. SC+CLM) could yield low accuracy on one task (SFT-CMS, 0.2) but high accuracy on others (SFT-RAG/SFT-CBQA, $\geq 0.6$).
@@ 4 Learning to Compare
4.1 Formulation
Given two LLMs $m_{i}$ and $m_{j}$, our goal is to predict which model achieves better downstream SFT performance. We denote the values of the five proxies for each model $m_{i}$ as $\{P_{m_{i}}^{k}\}_{k\in\mathcal{D}}$, where $\mathcal{D}=\{\text{PPL-CLM},\text{PPL-SC},\text{Kshot-CMS},\text{Kshot-RAG},\text{Kshot-CBQA}\}$. The learning-to-compare model leverages these proxies by training a binary classifier $f$ to predict the fine-tuned performance comparison between model pair $(m_{i},m_{j})$. For each proxy $k$, we construct the feature vector: $h_{k}(p_{m_{i}},p_{m_{j}})=\left[p^{k}_{m_{i}}-p^{k}_{m_{j}},\;p^{k}_{m_{i}}\cdot p^{k}_{m_{j}},\;p^{k}_{m_{i}},\;p^{k}_{m_{j}}\right]\in\mathbb{R}^{4}.$ We concatenate features from all five proxies to form the input and lead to 20 features, namely, $H(p_{m_{i}},p_{m_{j}})\in\mathbb{R}^{20}.$ We define the ground-truth label $y_{ij}$ as a binary value, where $y_{ij}=1$ if LLM $m_{i}$ performs better after SFT than $m_{j}$, and $y_{ij}=0$ otherwise. The classifier is trained by minimizing the binary cross-entropy loss (formulation is provided in Appendix Section D).
4.2 Experiment Setup
We implemented the supervised classifier using LightGBM (details on other models in Appendix Section D), training separate models per SFT task (CMS, RAG, CBQA). To ensure robustness, we performed 20 runs, each using a random 60%/40% split of the 50 LLM variants to generate training/testing pairs (splits varied per run). We report mean accuracy and standard deviation over the 20 runs in Table 2 (middle section), compared against unsupervised baselines including PPL-CLM and Kshot-RAG.
4.3 Results
@@
[…]
Learning To Compare (% Relative to Kshot-RAG)
@@ 5.2 Recall the Best Model from a Small Candidate Set
One key practical use for LLM performance predictors is to identify the most promising models within a group of candidates, which can lead to significant cost savings by reducing the number of models that undergo supervised fine-tuning. To assess our classifier’s effectiveness in this critical application—specifically, its ability to recall the best pre-trained LLMs—we performed a ranking experiment where pairwise comparisons between models were predicted and then aggregated into an overall ranking using Borda Count scoring Dwork et al. (2001). Specifically, for each model $m_{i}$, we compute its total score by counting the number of pairwise wins over all other models.
$\text{Score}(m_{i})=\sum_{j\neq i}\mathbf{1}_{\left(f(m_{i},m_{j})>0.5\right)}$
where $f(m_{i},m_{j})$ denotes the classifier’s predicted probability that $m_{i}$ outperforms $m_{j}$. $\mathbf{1}_{(\cdot)}$ is the indicator function. Finally, models are ranked based on their total scores, with higher scores indicating better predicted fine-tuned performance. Models achieving more pairwise ’wins’ received higher scores, indicating better predicted performance. The evaluation results, presented as top-1 and top-5 recall in Figure 5, show that our “learning-to-compare” method consistently identified the top-performing LLMs. Impressively, it achieved perfect top-1 recall for the SFT-CMS, SFT-RAG, and SFT-CBQA tasks by focusing on the top 7, 7, and 8 predicted models respectively, demonstrating its effectiveness even when narrowing down a relatively small candidate pool (as few as 8 models). Additionally, the unsupervised Kshot-RAG method showed strong performance, corroborating observations from Section 3.
@@ 6 Related Work
Pre-training of LLMs
LLM pre-training fundamentally shapes capabilities like reasoning (Wei et al., 2022; Kojima et al., 2022; Zellers et al., 2019), knowledge (Chang et al., 2024), and tool use (Yao et al., 2023; Mo et al., 2023). Critical pre-training design choices include the training objective—such as dominant CLM (Brown et al., 2020; OpenAI, 2023) for generation, SC (Raffel et al., 2020) which aids fine-tuning (Tay et al., 2023a), or combined UL2-style approaches (Tay et al., 2023a, b; Garcia et al., 2023), potentially using PrefixLM (Du et al., 2022; Chowdhery et al., 2023)—and pre-trained corpus composition, which involves quality curation (Rae et al., 2021; Touvron et al., 2023), filtering (Penedo et al., 2023; Xia et al., 2024), and source mixing (Weber et al., 2024; Shen et al., 2024) to ensure broad coverage and robustness. Given the variety of design options, lightweight methods to predict final performance are highly desirable for efficient model development. This work investigates predictors for supervised fine-tuning outcomes, utilizing systematic variations across several pre-training design factors in our study.
LLMs SFT Performance Prediction
The ability to predict the performance of large language models (LLMs) after fine-tuning has gained significant importance, largely driven by the substantial computational investment required for pre-training. Previous research (Kaplan et al., 2020; Hoffmann et al., 2022; Henighan et al., 2020) established scaling laws showing that increasing pre-training FLOPs typically reduces perplexity on held-out data, correlating with enhancements in capabilities like chain-of-thought reasoning (Wei et al., 2022; Kojima et al., 2022), preference alignment (Ouyang et al., 2022; Bai et al., 2022), and multilingual understanding (Chowdhery et al., 2023), suggesting larger models generally yield better downstream performance. Analogous scaling phenomena, where lower perplexity often corresponds to improved outcomes, have also been noted when fine-tuning LLMs for specific applications (Zhang et al., 2024; Isik et al., 2025); for instance, Isik et al. (2025) reported such a correlation for machine translation performance. However, token-level perplexity can over-weight frequent tokens and mask deficits on rare or semantically critical ones (Sinha and et al., 2020), and its predictability has been questioned for long-context generation (Liu et al., 2024b; Fang et al., 2025) and many-shot in-context learning (Agarwal et al., 2024), implying it may not be a robust indicator across all downstream tasks.
Departing from scaling-law studies across varying model sizes or from settings focused on extreme input/output lengths, we evaluate the efficacy of perplexity as a predictor of fine-tuned performance among same-size LLMs trained with the same pre-training compute, on widely used NLP tasks. In this controlled regime, we find that perplexity is not a reliable predictor of downstream SFT performance, calling into question its utility as a one-size-fits-all proxy for selecting among equal-size LLM variants. Building on this observation, we introduce several pre-training accessible proxies that exhibit stronger correlations with downstream SFT outcomes. And further propose a learning-to-compare framework that ensembles these proxies to rank candidate models, yielding consistent gains over any single proxy and outperforming perplexity-based selection.
@@ 7 Conclusion and Future Directions
Future work should explore the generalizability of these findings to larger model scales and a broader range of downstream tasks and fine-tuning paradigms. Further investigation into a broader array of pre-training strategies, data compositions, and the development of even more sophisticated proxy metrics could yield deeper insights. Additionally, exploring the theoretical connections between specific pre-training objectives or data characteristics and their influence on downstream task adaptability after fine-tuning represents a promising avenue for future work, ultimately enabling more efficient LLM development and selection.
@@ Acknowledgements
Appendix A LLM Usage Statement
We employed large language models (LLMs) mainly for refining the authors’ original writing, aiming to improve clarity and readability.
Appendix B Pretraining and LLMs
@@ 35 UL2R+CLM DC-0 1e-3
Appendix C Proxy Predictive Accuracy
@@
Appendix D Classifier Implementation Detail
@@
The implementation details are as follows: For logistic regression, we use scikit-learn’s (Pedregosa et al., 2011) LogisticRegression with the default lbfgs solver for binary classification. The model applies $L_{2}$ regularization with strength $C=1.0$, fits an intercept, and runs up to 100 iterations. Class weighting is not applied. For the neural network, we use scikit-learn’s MLPClassifier with two hidden layers of size 32 each and ReLU activation. The model is optimized using the Adam solver and trained for a maximum of 100 iterations. All other hyperparameters are set to their default values. For LightGBM, we use the LGBMClassifie from the official lightgbm library  [^1] https://lightgbm.readthedocs.io/en/latest/pythonapi/lightgbm.LGBMClassifier.html. The objective is set to binary with binary_logloss as the evaluation metric. All other hyperparameters follow the default settings: num_leaves=31, learning_rate=0.1, n_estimators=100, feature_fraction=1.0, bagging_fraction=1.0, and no regularization (lambda_l1=0.0, lambda_l2=0.0).